David Fifield

16816 E. Kenyon Dr. Aurora, CO 80013 (303) 766–3525

To the King Soopers Store Manager:

I am a Boy Scout working on my Eagle Scout Service Project. This is a project that all would-be Eagle Scouts must do, and it must benefit the community in a meaningful way. For my project, I have chosen to assist the Aurora Public Library. My project has two parts: first, a campaign to collect donations of books and money from the community; and second, a pancake breakfast fundraiser to raise money so that the library may purchase new books. It is with this second part of my project that I ask your assistance.

In order for the fundraiser to be successful, I must rely on corporate donations for things like paper plates, napkins, and plastic silverware, as well as the ingredients for the preparation of food. From having spoken to my fellow Scouts, I know that yours is a store that is generous towards causes like mine. With this in mind, I humbly ask you to donate any of the items mentioned above, or something equivalent such as a gift certificate. Remember that your donation not only helps me to achieve the rank of Eagle, it also benefits the Aurora Public Library.

If you have any questions, you may contact me by telephone at (303) 766–3525 or by

email at david@171recon.freeservers.com.

Sincerely yours,

David Fifield